The sum of two numbers is 27. The one number is 5 less than 3 times the other number. Find the numbers. Write and solve a system of equations to find the numbers
Answer by mananth(16946)


The sum of two numbers is 27.
Let one number be x
and other be y (y>x)
x+y=27----------------------(1)
The one number is 5 less than 3 times the other number.
y= 3x-5
substitute y in (1)
x+3x-5=27
4x-5=27
4x = 32
x =8
if x = 8 then 8+y=27
y = 19
the numbersare 19 & 8